# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v1

@@ 4 Numerical results
MVAR(4) 0.3599 (0.2408, 0.4535) 0.2536 (0.1609, 0.3098) 0.2251 (0.1361, 0.3435)
MVAR(9) 0.3783 (0.2865, 0.4655) 0.2765 (0.2040, 0.3367) 0.2569 (0.1935, 0.3330)
LSTM(4) 0.6323 (0.4313, 0.7941) 0.4160 (0.2889, 0.5269) 0.3372 (0.2532, 0.5037)
LSTM(9) 0.4646 (0.3172, 0.6820) 0.3185 (0.2151, 0.4626) 0.3010 (0.1993, 0.4393)
As shown in Table 2, the prediction accuracy of the closed-loop time stepper with the use of MVAR models outperforms the one with LSTMs, demonstrating robust generalization with tighter 10-90 percentile ranges. In addition, while MVARs with shorter lag demonstrate slightly higher approximation accuracy, LSTMs significantly improve with higher lags, albeit remaining inferior to MVARs. We hereby note that, considering the baseline 5-10% $L_{2}$ relative error of POD reconstruction (see Fig. 3), the $L_{2}$ errors are 3-10 times higher for the end-to-end closed-loop framework, underscoring the compounding effect of autoregressive prediction.
To highlight the robust performance of the proposed closed-loop time-stepping framework over time, we further depict the relative $L_{2}$ and $L_{\infty}$ errors, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in time, across the multiple cases of different, unseen initial conditions in Figs. 4 and 5 for MVARs and LSTMs, respectively. In particular, Figs. 4,5 show the average relative errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10-90 percentile ranges, over the $c=10$ cases considered, across time steps $k=1,\ldots,1100$.
@@
Figure 4: End-to-end numerical accuracy of the closed-loop framework across the testing set over time, using the trained MVAR models. The relative $L_{2}$ (red) and $L_{\infty}$ (green) errors of the high-dimensional density space, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (36) are displayed for the MVAR(4) (panel a) and MVAR(9) (panel b) latent dynamics models; average (solid) and 10-90 percentiles (dashed) are shown overall $C=10$ cases per time step.
Figures 4a,b) depict the results obtained with the MVAR(4) and MVAR(9) models, respectively, both indicating that, in spite of the long-time prediction, the relative error remains bounded in values below 30%; 5-10% is the baseline error of POD reconstruction in Fig. 3. In addition, the tight 10-90 percentile ranges indicate good generalization. Although both models exhibit comparable error profiles, the MVAR(4) model demonstrates slightly improved accuracy while using fewer parameters, as also reported in Table 2.
Similar bounded results are reported in Fig. 5 (a,b), depicting the employment of LSTM(4) and LSTM(9) models for long-time predictions.
(a) Relative Errors with LSTM(4)
(b) Relative Errors with LSTM(9)
Figure 5: End-to-end numerical accuracy of the closed-loop framework across the testing set over time, using the trained LSTM models. The relative $L_{2}$ (red) and $L_{\infty}$ (green) errors of the high-dimensional density space, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (36) are displayed for the LSTM(4) (panel a) and LSTM(9) (panel b) latent dynamics models; average (solid) and 10-90 percentiles (dashed) are shown overall $C=10$ cases per time step.
As expected by the results in Table 2, the use of LSTMs provides less accurate predictions in comparison to MVARs. However, LSTMs provide tighter 10-90 percentile ranges towards the end of time integration, indicating a more robust behavior in comparison to MVARs. In addition, as also reported in Table 2, a wider lag window makes the scheme with LSTMs more accurate, while that of MVARs less accurate. Nevertheless, when comparing across model classes, the MVAR(4) model achieves slightly lower relative errors overall, while also requiring significantly fewer trainable parameters and hyperparameters. This makes MVAR(4) not only more accurate in this setting but also more computationally efficient.
@@
For completeness, we provide a comparison of the “ground truth” normalized density fields $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ derived based on the closed-loop time-stepper in Eq. (35) for a particular case, in the testing data set, of an initially spread pedestrian group; Gaussian initialization corresponding to the 6th row of Table B.2. Figure 6 displays such a comparison for the MVAR(4) model in four selected time steps (k=250, 500, 780, 930k=250,\ 500,\ 780,\ 930) where the pedestrian crowd does not pass through the corridor boundary; similar comparisons are provided for the MVAR(9), LSTM(4) and LSTM(9) models in Figs. E.1, E.2 and E.3 of Appendix E, respectively.
@@
Figure 6: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq. (35) using the MVAR(4) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
As shown in Fig. 6, the predicted density fields preserve the global structure and propagation direction observed in the “ground truth” density, demonstrating the framework’s ability to capture dominant crowd dynamics and to deliver meaningful long-horizon predictions. The predicted density fields additionally succeed in correctly respecting the obstacle region. Comparison with the other (less accurate) models in Figs. E.1, E.2, E.3 reveals that while the above two crucial for crowd flow characteristics are preserved, the LSTM(4) model (the least accurate one) tends to underestimate peak density values, introduce artificial density in low density regions (for coping with conservation of mass), and exhibit greater dissipation in regions of high density; these erroneous behaviors relax as more accurate models are used, i.e., the LSTM(9) and the MVAR(9) ones.
@@
As shown, the proposed framework provides a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics. In particular, for the benchmark case of $100$ pedestrians moving in the corridor for more than 275s275\ $\text${s}, the conventional numerical integration requires a computational time of 20s20\ $\text${s} for the microscopic SFM simulator simulation plus 30s30\ $\text${s} for the extraction of the density fields (in a coarser time discretization), summing up to a total of 50s50\ $\text${s}. In contrast, our framework requires just $0.0063$, 0.0161s0.0161\ $\text${s} for the MVARs models and $6.75$, 6.77s6.77\ $\text${s} for the LSTM ones to complete the predictions across the long-time horizon integration–representing a speed-up of 3,100x and 7.4x for the MVARs and LSTMs, respectively. Even when accounting for the offline costs (POD basis construction and ROM training), MVARs retain a significant advantage, with training times under 0.1s0.1\ $\text${s} (as shown in Table 1). The MVARs variants prove particularly efficient for real-time applications, while maintaining better accuracy than their LSTM counterparts. These gains are especially valuable for real-world scenarios, where the number of pedestrians renders the conventional approaches unfeasible.
5 Conclusion
@@
In this work, we present a four-step data-driven framework that combines manifold learning and machine learning to construct a discrete evolution operator for emergent crowd dynamics. Our framework provides a highly computationally efficient and systematic approach for learning accurate ROMs for spatio-temporal emergent crowd dynamics. By first encoding the full high‐dimensional state into a compact latent representation, we learn the system’s evolution operator using surrogate ROMs—sidestepping the prohibitive computational cost of learning PDEs via for example convolutional neural networks (CNNs) or other high‐dimensional ML models such as DNNs. Working in latent space simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics.
@@
Thus, our approach does not attempt to learn any closed‑form PDE, but instead learns the effective solution operator of such an unavailable PDE, mapping initial crowd density to its future evolution. The resulting ROMs trained in the latent space, in the form of linear (MVARs) or nonlinear multivariate autoregressive models (such as LSTMs), encode and reconstruct this operator numerically via the POD reconstruction. As a result, our framework allows for rapid and accurate “what‐if” scenarios exploration and real‐time forecasting of crowd dynamics without sacrificing accuracy or stability, making them an ideal tool for large‐scale, real-world crowd‐dynamics analysis and control.
Finally, we note that linear MVARs exhibit a better performance in terms of the median relative $L_{1}$, $_{2}, LL_{\infty}$ errors when compared to LSTMs, thus offering a significantly lower complexity in their training and greater interpretability. This is in line with the results presented in other works [56, 94, 63, 57] which utilize PCA-based delay-coordinate embeddings to construct linear ROMs using in latent spaces to approximate complex nonlinear or chaotic dynamics.
@@ Appendix A Social Force Model
$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{i,j}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{i,B}(\bm{x}_{i})}_{\text{repulsive force}},$ (A.1)
where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is their maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time. The social force term in Eq. (A.1) introduces a short-range repulsive interaction between pedestrians, to prevent collisions. The social force between pedestrians $i$ and $j$ is given by:
@@
where $A_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction, $d_{ij}=\|\bm{x}_{i}-\bm{x}_{j}\|$ is the distance between pedestrians, $r_{i,j}\in\mathbb{R}$ is the sum of their radii (minimal comfortable distance) [64], $B\in\mathbb{R}^{+}$ is the parameter controlling the spatial decay rate of the repulsive force, $\bm{n}_{ij}=\big{(}n^{1}_{i,j},n^{2}_{i,j}\big{)}=\frac{\bm{x}_{i}-\bm{x}_{j}}{d_{ij}}$ is the unit vector pointing from pedestrian $j$ to pedestrian $i$, $g(x)$ is a function which ensures that physical forces are only active when pedestrians overlap, i.e., when $(d_{i,j}<r_{i,j})$. Moreover, the constant $k>0$ governs the strength of the linear contact (body) force, which causes repulsion between pedestrians when they are in contact. The third term is a sliding friction term that involves a coefficient $\kappa>0$, which scales with the tangential component of the relative velocity $\Delta v_{ij}^{\mathrm{tang}}$; this component is projected onto the unit tangential vector $\mathbf{t}_{i,j}=\big{(}-n^{2}_{i,j},n^{1}_{i,j}\big{)}$. The repulsive force in Eq. (A.1), accounting for the repulsion that a pedestrian might experience due to static objects $B$, follows a similar, to Eq. (A.2), form, reading:
$\mathbf{F}_{iB}(\bm{x}_{i})=C_{i}\exp\left(-\frac{d_{iB}}{D_{i}}\right)\bm{n}_{iB}+k\,g(r_{i}-d_{iB})\,\bm{n}_{iB}.$ (A.3)
@@
Here, $C_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction against a wall or an obstacle and $D_{i}$ is its corresponding decay parameter. Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exert dynamic resistance.
For our simulations, the model parameters were calibrated based on established values from previous studies [1, 64]. These works describe pedestrians with a standard average mass of 8080kg, for which a relaxation time $\tau=0.5\,\text{s}$ provides adequate behavior under normal conditions. The complete set of parameters governing repulsive potentials, including agent-agent and agent-obstacle interactions, along with the simulation environment specifications, are detailed in Table A.1. This parameter selection, grounded in empirical crowd studies, enables the simulation of diverse scenarios, including the present case study while maintaining consistency with validated bottleneck evacuations scenarios, and even though the parameters should vary individually, in the literature the parameters are often chosen equally for all pedestrians to maintain a reasonable number of parameters calibration [64].
Table A.1: Simulation parameters.
Simulation Parameters Value
Number of Pedestrians 100
Corridor length ($L_{x}$) $48\,\text{m}$
Corridor width ($L_{y}$) $12\,\text{m}$
Time Step $\delta_{t}$ $2.5\times 10^{-2}\,\text{s}$
Simulation time $275\,\text{s}$
@@
Repulsion coefficient ($k$) $1.2\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$

### v2

@@ 4 Numerical results
MVAR(4) 0.127(0.072,0.201)0.127\ (0.072,0.201) 0.078(0.045,0.119)0.078\ (0.045,0.119) 0.073(0.039,0.117)0.073\ (0.039,0.117)
MVAR(9) 0.127(0.072,0.200)0.127\ (0.072,0.200) 0.078(0.045,0.118)0.078\ (0.045,0.118) 0.072(0.039,0.116)0.072\ (0.039,0.116)
LSTM(4) 0.280(0.265,0.325)0.280\ (0.265,0.325) 0.167(0.158,0.190)0.167\ (0.158,0.190) 0.138(0.121,0.167)0.138\ (0.121,0.167)
LSTM(9) 0.309(0.297,0.351)0.309\ (0.297,0.351) 0.185(0.174,0.207)0.185\ (0.174,0.207) 0.135(0.115,0.167)0.135\ (0.115,0.167)
As shown in Table 2, the open-loop (one-step) prediction accuracy with the use of MVAR models outperforms that of LSTMs; all LSTM relative errors are approximately twice as high as those of the MVAR models. However, LSTMs demonstrate narrower 10–90% error percentiles in comparison to MVARs, indicating less uncertainty in one-step predictions. Table 2 additionally shows that the accuracy of MVARs does not improve with a longer lag, while that of LSTMs slightly degrades.
@@
(c) Relative Errors with LSTM(4)
(d) Relative Errors with LSTM(9)
@@
Figure 4: One-step (open-loop) prediction errors, in the ambient–density profile–space, for the testing set over time, using trained MVAR and LSTM models. The relative reconstructed $L_{2}$ (red) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ Eq. (37) (using the one-step prediction in Eq. (35)) are shown for the MVAR(4) panel (a), MVAR(9) panel (b), LSTM(4) panel (c), and LSTM(9) panel (d) latent dynamics models. Mean relative errors (solid) and 10–90% error percentiles (dashed) are shown over $C=20$ cases per time step. For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
Fig. 4 displays the average relative reconstructed errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10–90% percentile ranges, for one-step predictions, over the $C=20$ cases of different, unseen initial conditions considered, across the time steps $k=1,\ldots,1000$. Fig. 4(a,b) depict the results obtained with the MVAR(4) and MVAR(9) models, respectively, both demonstrating a remarkable approximation accuracy. The mean $L_{2}$ reconstructed relative errors are initially $\sim$15% and reduce to $\sim$7% after the first 50s50\ $\text${s} of the simulation. This level of errors matches the infimum approximation error baseline of POD reconstruction, shown in Fig. 3(b). This very high approximation accuracy is not met by the LSTM(4) and LSTM(9) models, shown in Fig. 4(c,d). In particular, the mean relative reconstructed $L_{2}$ errors of the LSTM(4) and LSTM(9) models are $\sim$17% and $\sim$18%, respectively, towards the end of the simulations. Fig. 4 further confirms that the lag window does not affect the one-step prediction accuracy of MVARs, while it slightly reduces that of LSTMs.
Recursive/Closed-loop Predictions
For obtaining long-term (closed-loop) predictions with the proposed framework, we next adopt the recursive closed-loop time-stepper in Eq. (36), which is only supplied with the first $w$ observations and predicts the next time steps in an autoregressive fashion. To quantify the prediction accuracy, we again use the relative reconstructed $L_{1}$, $L_{2}$, and $L_{\infty}$ errors in the ambient space, computed for the cases included in the testing set $X_{ts}$ across time steps. Table 3 provides a summary of these reconstructed errors, including the mean and the 10–90% error percentiles, aggregated over all cases and time steps.
Table 3: Closed-loop/recursive simulation errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models. The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (37) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models. Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
MVAR(4) 0.231(0.139,0.376)0.231\ (0.139,0.376) 0.153(0.087,0.267)0.153\ (0.087,0.267) 0.141(0.074,0.247)0.141\ (0.074,0.247)
MVAR(9) 0.184(0.114,0.292)0.184\ (0.114,0.292) 0.124(0.073,0.209)0.124\ (0.073,0.209) 0.119(0.062,0.210)0.119\ (0.062,0.210)
LSTM(4) 0.419(0.221,0.648)0.419\ (0.221,0.648) 0.284(0.139,0.512)0.284\ (0.139,0.512) 0.281(0.115,0.501)0.281\ (0.115,0.501)
LSTM(9) 0.433(0.230,0.590)0.433\ (0.230,0.590) 0.310(0.145,0.460)0.310\ (0.145,0.460) 0.285(0.121,0.441)0.285\ (0.121,0.441)
As shown in Table 3, the long-term (closed-loop) prediction accuracy of the proposed framework with MVAR models outperforms significantly that of LSTM models. In addition, MVARs demonstrate narrower 10–90% error percentiles than LSTMs, indicating more reliable long-term predictions. Table 3 further shows that the consideration of a longer lag makes MVARs more accurate, while slightly degrading the accuracy of LSTMs.
We next assessed the evolution/accumulation of the relative reconstructed errors $L_{2}$ and $L_{\infty}$ over time for the closed-loop setting. Fig. 5 depicts the average reconstructed relative errors in the ambient–density profile–space, and their 10–90% error percentile ranges across the multiple cases of different, unseen initial conditions in the testing set for all four ROMs.
(a) Relative Errors with MVAR(4)
(b) Relative Errors with MVAR(9)
(c) Relative Errors with LSTM(4)
(d) Relative Errors with LSTM(9)
(e) Relative Errors with SFM, $1\%$ perturbation
(f) Relative Errors with SFM, $1\%$ perturbation
Figure 5: Closed-loop/recursive simulation errors in the ambient–density profile–space, across the testing set over time, using the trained MVAR and LSTM models. Panels (a)–(d) display the relative $L_{2}$ (red) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ in Eq. (37) for MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models, respectively. Panels (e) and (f) provide a system’s “baseline” relative error estimation, obtained from the SFM simulations using $\pm 1\%$ perturbations in the initial conditions for the density profiles (see in Section 3.1). Mean relative error (solid) and 10–90% percentiles (dashed) are shown over $C=20$ cases per time step. For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
Indicatively, for our illustrations, to provide a comparative “baseline” of the reconstructed error for tiny perturbations of the density profiles, we quantified the sensitivity of the SFM simulations per se with respect to small perturbations of the initial conditions (see in Section 3.1). Specifically, we introduced $\pm 1\%$ perturbations in the initial conditions of the density profiles of the reference test cases and computed the relative reconstructed errors between these perturbed initial conditions and their unperturbed counterparts. The MVAR models Fig. 5(a,b) demonstrate a remarkable approximation accuracy, with the mean relative $L_{2}$ error remaining bounded below $\sim$14%(17%) for the MVAR(9)(MVAR(4)) models, despite the long-term horizon. The respective 10–90% error percentiles are bounded in (10, 30)% relative errors for the MVAR(4) model, while for the MVAR(9) model, the 10–90% error percentiles reduce further to (7, 21)% relative errors. The mean values of errors, and especially of the MVAR(9) model, is well within the 10–90% error percentiles of the system’s deviation/“baseline” 10–90% $L_{2}$ error percentiles (2, 18)% relative error, as quantified by the SFM simulations of the testing set using the $\pm$ 1% perturbed initial conditions for the density profiles, shown in Fig. 5(e,f). In contrast, the LSTM models Fig. 5(c,d) provide less accurate predictions than the MVAR models, with mean relative $L_{2}$ error plateauing at $\sim$35% for both LSTM(4) and LSTM(9) models, towards the end of the simulation, well above the expected “baseline” error bounds.
@@
For completeness, we directly compare the ground truth normalized density fields $x^{(c)}(t_{k})$ and the predicted fields $\hat{x}^{(c)}(t_{k})$ generated by the closed-loop/recursive simulations. This analysis uses a case from the test dataset featuring a Gaussian initialization corresponding to the 6th row of Table B.2. Figure 6 displays such comparison for the MVAR(9) model at four selected time steps (k=250, 500, 780, 930k=250,\ 500,\ 780,\ 930), during which the pedestrian crowd does not pass through the corridor boundary. Panels (a,c,e,g) show the ground truth at these time steps, while panels (b,d,f,h) present the corresponding MVAR(9) model predictions; analogous results for the MVAR(4), LSTM(4) and LSTM(9) models are provided in Appendix E, Figs. E.1, E.2 and E.3, respectively.
@@
Figure 6: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (36) using the MVAR(9) model for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
The predictions from all four surrogate models successfully preserve the global structure and propagation direction of the ground truth density, demonstrating the framework’s capacity to capture dominant crowd dynamics and produce meaningful long-horizon predictions. The models also correctly respect the obstacle regions. The primary observed discrepancy is a minor underestimation of the peak density values coupled with the introduction of artificial density in low-density regions, a likely consequence of enforcing mass conservation via the POD restriction operator. This artifact is faintest in the MVAR(9) model, and remains only slightly more pronounced in the MVAR(4). By contrast, it becomes considerably stronger in the less accurate LSTM(4) and LSTM(9) models.
MVAR vs. LSTM performance
The comparison between model classes reveals that the use of MVARs provides a more accurate and reliable framework for long-term predictions than that of LSTMs. This is clearly demonstrated by their significantly lower errors and narrower error percentiles in the closed-loop setting (Table 3, Fig. 5). Crucially, this performance gap widens substantially when compared to the open-loop results (Table 2, Fig. 4), where LSTMs performed more competitively. For instance, the MVAR(9) model’s mean $L_{2}$ error across time increased from $\sim$7% (open-loop) to $\sim$14% (closed-loop) within the “baseline” reconstruction error from POD and perturbed initial conditions, respectively, while the LSTM(9) model’s error changing respectively from $\sim$19% to $\sim$35% is well above these “baselines”. This superior performance of MVARs is attributed to the challenge of error accumulation in closed-loop forecasting; the linear structure of the MVAR model proves to be much less susceptible to this instability than complex, nonlinear LSTMs. While LSTMs can theoretically capture richer nonlinear dynamics, their performance severely degrades due to the “curse of dimensionality” in their training, i.e., finding a global optimum, accompanied by a distribution shift [99, 100] (i.e., the error accumulation of the recursive scheme during inference, that shifts the distribution upon which the models were trained), despite their ability to learn the one-step dynamics effectively. In addition, there are several challenges regarding the “correct” initialization of LSTMs [101], especially in dynamical systems and multiscale modeling tasks. In contrast, the training of the linear MVAR models is based only on OLS, thus giving a global optimum. For the particular applications, the results suggest that MVARs are a much better choice for long-term horizon predictions in the latent space. This is a result in line also with other studies (see e.g., [62]).
@@
Beyond accuracy, the proposed framework offers a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics. In particular, for the benchmark case of $100$ pedestrians moving in the corridor for $250\text{s}$, the conventional numerical integration requires a computational time of $20\text{s}$ for the microscopic SFM simulator simulation plus $30\text{s}$ for the extraction of the density fields (in a coarser time discretization), summing up to a total of $50\text{s}$. In contrast, our framework requires just $0063\text{s}, 0.0$$0161\text{s}$ for the MVARs models and $6.75\text{s}$, $6.77\text{s}$ for the LSTM ones to complete the predictions across the long-time horizon integration–representing a speed-up of 3,100x and 7.4x for the MVARs and LSTMs, respectively. Even when accounting for the offline costs (POD basis construction and ROM training), MVARs retain a significant advantage, with training times under $0.005\text{s}$ (as shown in Table 1). The MVARs variants prove particularly efficient for real-time applications, while maintaining better accuracy than their LSTM counterparts. These gains are especially valuable for real-world scenarios, where the number of pedestrians renders the conventional approaches infeasible.
5 Conclusions
@@
In this work, we present a four-step data-driven framework that combines manifold and machine learning to construct a discrete evolution operator for emergent crowd dynamics. Our framework provides a computationally efficient and systematic approach for learning ROMs for spatio-temporal emergent crowd dynamics. By first encoding the full high‐dimensional state into a compact latent representation, we learn the system’s evolution operator using surrogate ROMs, bypassing the prohibitive computational cost of learning PDEs via, for example, CNNs or other high‐dimensional ML models such as DNNs. Working in latent space with delayed POD coordinates simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics. In fact, under the usual generic assumptions of Takens’ theorem, applying time-delay embeddings to POD modes produces a delay-coordinate latent model that is diffeomorphic to the original system dynamics (see also results and discussion in [67, 62, 102, 68, 69]).
@@
Thus, our approach does not attempt to learn any closed‑form PDE, but instead learns the effective solution operator of such an unavailable PDE, mapping initial crowd density to its future evolution. The resulting ROMs trained in the latent space, in the form of linear (MVARs) or nonlinear multivariate autoregressive models (such as LSTMs), encode and reconstruct this operator numerically via the POD reconstruction. As a result, our framework allows for rapid and accurate “what‐if” scenarios exploration and real‐time forecasting of crowd dynamics. Critically, we demonstrate that linear MVAR models within our framework not only outperform more complex LSTMs in long-term predictive accuracy and reliability, but also achieve a computational speed-up of over three orders of magnitude compared to conventional simulation. This combination of interpretability, speed, and accuracy makes our approach, particularly with MVARs, an efficient approach for large‐scale, real-world crowd‐dynamics analysis and distributed control [103, 104, 19, 21, 24, 22, 105].
Finally, our finding that simple linear MVAR models outperform nonlinear LSTM networks in the latent space in terms of complexity, training time, and interpretability is a significant insight. It aligns with a growing body of literature [62, 102, 69, 63] that demonstrates the power of linear ROMs constructed in judiciously chosen latent spaces (e.g., POD-based delay-coordinate embeddings, Koopman operator) for approximating complex nonlinear or chaotic dynamics, offering a promising path toward interpretable and efficient data-driven modeling.
Acknowledgments
C.S. acknowledges partial support by the PNRR MUR projects PE0000013-Future Artificial Intelligence Research-FAIR & CN0000013 CN HPC - National Centre for HPC, Big Data and Quantum Computing, Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM). I.K. acknowledges partially support by the US National Science Foundation.
@@ Appendix A Social Force Model
$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{ij}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{iB}(\bm{x}_{i})}_{\text{repulsive force}},$ (A.1)
where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is the maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time.
The social force term in Eq. (A.1) introduces a short-range repulsive interaction between pedestrians to maintain social distances and avoid physical contact [1]. Its expression between two pedestrians $i$ and $j$ is given by:
@@
where the first term represents a repulsive force modelling the physiological desire for personal space, the second term represents a physical contact (body) force that is only active when pedestrians are in contact, and the third term represents a sliding friction force that occurs during physical contact. The repulsive force is modelled in Eq. (A.2) by an exponential term, where $A_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction, $d_{ij}=\|\bm{x}_{i}-\bm{x}_{j}\|_{2}$ is the euclidean distance between pedestrians, $r_{ij}=r_{i}+r_{j}\in\mathbb{R}$ is the sum of their radii (minimal comfortable distance [70]), $B_{i}\in\mathbb{R}^{+}$ is the parameter controlling the spatial decay rate of the repulsive force, and $\bm{n}_{ij}=\big(n^{1}_{ij},n^{2}_{ij}\big)=(\bm{x}_{i}-\bm{x}_{j})/d_{ij}$ is the unit vector pointing from pedestrian $j$ to pedestrian $i$. The physical contact force is modelled in Eq. (A.2) by a linear term, where $k>0$ controls the strength of the contact (body) force, and $g(x)=\text{max}(0,x)$ is a ramp function ensuring that physical forces are only active when pedestrians overlap, i.e., when $(d_{ij}<r_{ij})$. The friction force term in Eq. (A.2) is modelled by the same ramp function $g(x)$, scaled by a friction coefficient $\kappa>0$ and difference in the tangential velocities $\Delta v_{ij}^{\mathrm{tang}}=\|(\bm{v}_{i}-\bm{v}_{j})-\left((\bm{v}_{i}-\bm{v}_{j})\cdot\bm{n}_{ij}\right)\bm{n}_{ij}\|_{2}$; this component is projected onto the unit tangential vector $\mathbf{t}_{ij}=\big(-n^{2}_{ij},n^{1}_{ij}\big)$.
The repulsive force in Eq. (A.1), accounting for the repulsion that a pedestrian might experience due to static objects $B$, follows a similar form to Eq. (A.2), reading:
$\mathbf{F}_{iB}(\bm{x}_{i})=C_{i}\exp\left(-\frac{d_{iB}}{D_{i}}\right)\bm{n}_{iB}+k\,g(r_{i}-d_{iB})\,\bm{n}_{iB},$ (A.3)
@@
where $C_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction against a wall or an obstacle, $D_{i}$ is its corresponding decay parameter, and $d_{iB}$ is the euclidean distance between the pedestrian and the closest point of the object $B$; $\bm{n}_{iB}$ is the respective the unit vector. Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exerft dynamic resistance.
Table A.1: Social Force Model simulation parameters.
Simulation Parameters Value Simulation Parameters Value
@@
Number of Pedestrians ($N$) 100 Repulsion coefficient ($k$) $1.2\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$